OpenStudy (anonymous):

The equation of line AB is y = 5x + 1. Write an equation of a line parallel to line AB in slope-intercept form that contains point (4, 5). y = 5x − 15 y = 5x + 15 y = 1 over 5x + 21 over 5 y = 1 over 5x − 29 over 5

OpenStudy (dan815):

Do you know what the slope of y=5x+1 is?

OpenStudy (anonymous):

@dan815 i dont:?

OpenStudy (dan815):

y=mx+b <--- equation of a line m is slope, b is y-intercept

OpenStudy (dan815):

so what do you think the slope of y=5x+1 is

OpenStudy (anonymous):

Im sorry i have no idea im so lost);

OpenStudy (dan815):

you see how the m is the slope and its right before the x

OpenStudy (dan815):

similarly in y=mx+b y=5x+1 <----- the 5 is in the place of m, so the 5 is the slope

OpenStudy (dan815):

ok?

OpenStudy (anonymous):

the slope on a parallel line is the same so m=m but the y-intercept is different so b on the 1st equation is different from the second basic equation being: y=mx+b

OpenStudy (dan815):



OpenStudy (anonymous):

dan you still marked the bad line as parallel

OpenStudy (anonymous):

so the equation you are looking for starts off y=5x figure out where b is compared to the slope and where it crosses the y-axis

OpenStudy (anonymous):

Do you know how to use point slope form?

OpenStudy (dan815):

no i follow my own notation that arrow in the middle is equivalent to putting 2 arrow heads on the ends of the line saying they are continued for all x

OpenStudy (anonymous):

To write an equation of a line parallel to line AB in slope-intercept form that contains point (4, 5) I would use point slope form. A line that is parallel will have the same slope. The line y = 5x + 1 has a slope of 5. Use the point (4,5) with the point slope form to find a line that is parallel.